OpenStudy (anonymous):

What is the value of n? 9 24 14 26

OpenStudy (anonymous):

lol. Where's that picture of satellite73's?

OpenStudy (anonymous):

sorry there it is

OpenStudy (anonymous):

All internal angles need to sum to 180. So how could you go about working it out?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

add them all then equal it to 180?

OpenStudy (anonymous):

Yeah pretty much. (5n + 5) + 40 + 15 = 180 And work from there.

OpenStudy (anonymous):

jst solve by this 55+5n+5=180

OpenStudy (anonymous):

i got it thank you

OpenStudy (anonymous):

so n= 120 /5